OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. –2x – y = –7 2x – y = –7 2x – y = 11 –2x + y = –11

OpenStudy (anonymous):

2x-y=-7

OpenStudy (anonymous):

y-5=2(x+1) y-5=2x+2 -2x+y=7 2x-y=-7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

someonenamedmeme i nees help in another one

OpenStudy (anonymous):

y -5 = 2( x +1) y - 5 = 2x + 2 =>2x -y = -7